OpenStudy (anonymous):

7x+8y=5 -9x-2y=-23 x=? y=?

OpenStudy (anonymous):

Do you know substitution?

OpenStudy (anonymous):

Where you solve for x or y using one equation and plugging that into the other one.

OpenStudy (anonymous):

you can add them. first, you multiply 7x+8y=5 by 4 then you add both together 7x+8y= 5 -36x-8y=-92